# Diorama: Unleashing Zero-shot Single-view 3D Scene Modeling

### v1

@@ A.4 Multimodal 3D shape retrieval
For each observed object in the image, we use a two step process for retrieving matching shapes from our shape database. We first use a text query to retrieve objects matching the semantic category, and then image query to re-rank the retrieved candidates. We find this two-stage approach helps ensure that retrieved objects are of the correct semantic class, as it is possible for semantically different objects to be geometrically similar (e.g. books vs cardboard box).
@@
Given a pair of query image $I_{q}$ and reference rendering image $I_{r}$, we construct a “cyclical distance” map and define a similarity score using their corresponding DINOv2 patch features, $f_{q}$ and $f_{r}$, and feature masks $m_{q}$ and $m_{r}$. For each query feature $f_{q}^{i}$ at the patch location $i$, we compute its cyclical patch $i^{\prime}$ in $I_{q}$ as follows:
@@
where $S(\cdot,\cdot)$ denote as cosine similarity and a cyclical distance map is constructed as $D$ with $D_{i}=-||i,i^{\prime}||_{2}$. In the end, we build the feature correspondences by taking at most K query-reference patch pairs $(i,j)\in\mathcal{N}$ with top-K minimal cyclical distances in $D$. To ensure only keeping correspondences with strong similarity, we empirically set similarity score threshold as 0.7. The overall similarity between $I_{q}$ and $I_{r}$ is defined as the average of similarities of feature correspondences:
@@ A.6 Scene optimization
During scene optimization, we take the retrieved objects, initial object pose estimates, and optimize the placements so that the relationships in the scene graph is respected. In the main paper, we described the different stages of optimization. Here we provide more details on how support is enforced while ensuring there is sufficient space, and no inter-penetrations.
Support. A key relationship is that of support, for which we want to ensure that the each object is properly place in contact with their supporting object. To do so, for each object in our database, we precompute candidate support surfaces by identifying roughly planar surfaces on the object.
Based on estimated initial object poses and the scene graph $\mathbf{G}$ describing object relationships, for each a pair of interacting objects, we compute a pair of contact surface from the supported object and support surface from the supporting object. The support surface is determined by selecting the surface with the minimal distance to the centre of the supported object. The contact surface is selected to be the one having the closest direction with the support direction.
@@
Optimization. In each optimization stage, we use a separate SGD optimizer with initial learning rate 0.01 and momentum 0.9 for corresponding pose parameters, except for the Space stage where we set initial learning 0.001 for the scale parameter. We also decay the learning rate by 0.1 every 50 steps. We run 200 steps in total in each optimization stage.
@@ Appendix B Additional experiments
Implementation details. We render SSDB scene images of size 1008$\times$784. Considering computation efficiency and good coarse pose selection, we render 180 gray-color multiviews of size 224$\times$224 for each 3D shape from predefined camera viewpoints to focus on geometry-wise semantic similarity and leave out effect of texture. For zero-shot pose estimation, we use the fine-tuned ViT-L of DinoV2 [65] following [63] to embed image patches of size $14\times 14$.
B.1 Architecture reconstruction evaluation
@@
Metrics. For architecture reconstruction, we report 2D IoU, Pixel Error (PE), Edge Error (EE) and Root Mean Squared Error (RMSE) following [75]. Additionally, we introduce bounding-box-centered chamfer distance (CDb) to measure the correctness of 3D structure of the predicted architecture. Additionally, we measure runtime for each method, averaged per scene.
Results. We compare our proposed PlainRecon with RaC in Tab. 4 on the 344 images of SSDB scene, with predicted depth from DepthAnythingV2 (DAv2) and Metric3D (M3D). We find that RaC fails in roughly 30% of the cases, while PlainRecon is able to handle all inputs. This is because RaC sets up architecture reconstruction as a constrained discrete optimization problem, which can occasionally produce infeasible solutions resulting in no output. PlainRecon does not involve such an optimization and therefore avoids this shortcoming. M3D depth is overall superior to DAv2 when used by any of the methods. PlainRecon closely matches or outperforms RaC on most of the metrics, RMSE being an outlier where PlainRecon underperforms significantly. Notably, PlainRecon outperforms in terms of CDb meaning that it captures the overall structure of the architecture better than RaC. As RaC fails in a significant number of scenes, PlainRecon is clearer superior in terms of reliability.
B.2 Performance for different object groups
@@ B.3 Multiple retrievals
B.4 Different optimization strategy
@@
B.5 Quantitative results on ScanNet
Tab. 6 shows quantitative comparison between ROCA, DiffCAD and ours on the proposed evaluation set. We note that ROCA is trained end-to-end using imperfect Scan2CAD annotations and DiffCAD is trained on the synthetic data per category. Both ROCA and DiffCAD cannot generalize to unseen objects during training (indicated as gray-colored numbers). Our zero-shot method achieves competitive performance against DiffCAD that further exhibits performance degradation under the probabilistic setting due to the partial observations of commonly occluded objects. Fig. 13 visually shows differences between ground-truth evaluation sets used in DiffCAD, ROCA and ours.
@@ Appendix C Qualitative examples
We provide additional examples of generated scenes in Figures 14, 15 and 16. In Figure 14, we provide plausible arrangements based on renders from SSDB, as well as the associated predicted scene-graphs. With Diorama, we can produce alternative arrangments that uses different objects, while respecting the spatial relationships of the original image (e.g. picture on the wall, monitor on the desk).

### v2

@@ A.4 Multimodal 3D shape retrieval
For each observed object in the image, we use a two-step process for retrieving matching shapes from our shape database. We first use a text query to retrieve objects matching the semantic category, and then an image query to re-rank the retrieved candidates. We find this two-stage approach helps ensure that retrieved objects are of the correct semantic class, as it is possible for semantically different objects to be geometrically similar (e.g. books vs cardboard box).
@@
We pre-compute $T$ multiview renderings and depth maps for each retrieved CAD model $s_{i}$. Each query object crop $I^{o_{i}}$ and set of multiview images $\{I^{s_{i,j}}\}$ is encoded into normalized patch features $\mathcal{F}(I)$ using DinoV2. For each patch embedding from the query image, we construct a semantic correspondence to a patch embedding from each rendering view with the minimal cyclical distance [26, 56]. We compute correspondence score as cosine similarity between patch embeddings. Specifically, given a pair of query image $I_{q}$ and reference rendering image $I_{r}$, we construct a “cyclical distance” map and define a similarity score using their corresponding DINOv2 patch features, $f_{q}$ and $f_{r}$, and feature masks $m_{q}$ and $m_{r}$. For each query feature $f_{q}^{i}$ at the patch location $i$, we compute its cyclical patch $i^{\prime}$ in $I_{q}$ as follows:
@@
where $S(\cdot,\cdot)$ denote as cosine similarity and a cyclical distance map is constructed as $D$ with $D_{i}=-||i,i^{\prime}||_{2}$. In the end, we build the feature correspondences by taking at most K query-reference patch pairs $(i,j)\in\mathcal{N}$ with top-K minimal cyclical distances in $D$. To ensure only keeping correspondences with strong similarity, we empirically set similarity score threshold as 0.7. The overall similarity score between $I_{q}$ and $I_{r}$ is defined as the average of similarities of feature correspondences:
@@
The query image is matched to the most similar multiview rendering with the maximal averaged correspondence-wise feature similarity to produce a coarse pose hypothesis.
@@ A.6 Scene optimization
During scene optimization, we take the retrieved objects, initial object pose estimates, and optimize the placements so that the relationships in the scene graph are respected. In addition to support relations described in the scene graph, we further establish an adherence relationship between each object on the floor and potential walls using a heuristic algorithm that judging whether the closest distance from the object’s surface center to a candidate wall is within a predefined threshold. We hence also optimize object position and rotation based on both the support and adherence relationship. In the main paper, we described the different stages of optimization. Here we provide more details on how support is enforced while ensuring there is sufficient space, and no inter-penetrations.
Support. A key relationship is that of support, for which we want to ensure that each object is properly placed in contact with their supporting object. To do so, for each object in our database, we precompute candidate support surfaces by identifying roughly planar surfaces on the object. Based on estimated initial object poses and the scene graph $\mathbf{G}$ describing object relationships, for each a pair of interacting objects, we compute a pair of contact surface from the supported object and support surface from the supporting object. The support surface is determined by selecting the surface with the minimal distance to the center of the supported object. The contact surface is selected to be the one having the closest direction with the support direction.
@@
Optimization. In each optimization stage, we use a separate SGD optimizer with initial learning rate 0.01 and momentum 0.9 for corresponding pose parameters, except for the Space stage where we set initial learning 0.001 for the scale parameter. We also decay the learning rate by 0.1 every 50 steps. We run 200 steps in total in each optimization stage. We describe the objective function of each optimization stage as below, including weight hyperparameters:
Stage 1: $e_{1}=3\cdot e_{\text{align}}+e_{\text{sem}}$ (1)
Stage 2: $e_{2}=5\cdot e_{\text{place}}+e_{\text{rel}}$ (2)
Stage 3: $e_{3}=e_{\text{vol}}$ (3)
Stage 4: $e_{4}=5\cdot e_{\text{place}}+e_{\text{col}}$ (4)
@@ Appendix B Additional experiments
Implementation details. We render SSDB scene images of size 1008$\times$784. Considering computation efficiency and good coarse pose selection, we render 180 gray-color multiviews of size 224$\times$224 for each 3D shape from predefined camera viewpoints to focus on geometry-wise semantic similarity and leave out effect of texture. For zero-shot pose estimation, we use the fine-tuned ViT-L of DinoV2 [58] following [56] to embed $14\times 14$ image patches. We run experiments on one Nvidia RTX 4090 GPU.
B.1 Details of comparing against ACDC
We run ACDC on all SSDB images with the default configuration. For both ACDC and Diorama, we feed ground-truth 2D object bounding boxes and segmentation masks into systems to avoid 2D perception errors for analysis. Both systems retrieve object from out-of-distribution 3D shape collections. Since ACDC is developed upon the OmniGibson simulation platform, it retrieves from the built-in set of approximately 8,800 OmniGibson CAD objects for convenient deployment in the physics engine. For Diorama, since we do not specify certain CAD file formats to be compatible with a physics engine, we are able to include more 3D shapes for retrieval from different sources. In particular, we compose a set of 25K 3D shapes for Diorama to retrieve from. For ACDC, overall runtime is dominated by network API calls to the LLM (GPT4o). ACDC calls GPT4o three times per detected object, while Diorama calls GPT4o twice in total, irrespective of the number of objects. For the user study, we randomly sample 48 images and corresponding results to ask the participants to assess the quality of single-view 3D scene modeling in terms of object matching and overall scene quality. For object matching, we consider both semantic correctness and geometric similarity. For overall scene quality, we consider the accuracy of architecture reconstruction and object arrangement, and physical plausibility of the whole scene. The question order is randomly shuffled to the participant.
B.2 Architecture reconstruction evaluation
@@
As ACDC outputs plane parameters and masks but does not output actual mesh planes, we perform a simple extraction procedure. First, we back-project the depth used by ACDC to the point cloud. Then, similarly to our method, we run RANSAC for each architectural element based on image segmentation masks to fit the plane, resulting in selecting inlier points that correspond to the largest planar region of the point cloud. We obtain the oriented bounding box from the inliers and extract mesh from it. Finally, we align the normals of the architectural elements with the normals ACDC used to optimize object placement.
B.3 Object alignment on estimated depth
Tab. 6 presents the comparison of different zero-shot pose estimation methods on the 9D CAD alignment task given predicted depth by Metric3D. The results align with our observation under the ground-truth depth setting.
B.4 Performance for different object groups
@@ B.5 Multiple retrievals
B.6 Different optimization strategy
@@
B.7 Quantitative results on ScanNet
Tab. 10 shows quantitative comparison between ROCA, DiffCAD and ours on the proposed evaluation set. Following prior work [27, 42, 24], we report the object alignment accuracy where a CAD model is considered correctly aligned if the translation error $\leq 20$cm, the geodesic rotation error $\leq 20^{\circ}$, and the scale ratio $\leq 20\%$. We note that ROCA is trained end-to-end using imperfect Scan2CAD annotations and DiffCAD is trained on the synthetic data per category. Both ROCA and DiffCAD cannot generalize to unseen objects during training (indicated as gray-colored numbers). Our zero-shot method achieves competitive performance against DiffCAD that further exhibits performance degradation under the probabilistic setting due to the partial observations of commonly occluded objects. Fig. 14 visually shows differences between ground-truth evaluation sets used in DiffCAD, ROCA and ours.
@@ Appendix C Qualitative examples
We provide additional examples of generated scenes in Figures 15, 16, 17 and 18. In Figure 16, we provide plausible arrangements based on renders from SSDB, as well as the associated predicted scene-graphs. With Diorama, we can produce alternative arrangements that use different objects, while respecting the spatial relationships of the original image (e.g. picture on the wall, monitor on the desk).